# RoboMIND: Benchmark on Multi-embodiment Intelligence Normative Data for Robot Manipulation

### v2

@@ VII Limitations
One limitation of RoboMIND is the relatively simple background environments. While our primary focus has been on constructing a large-scale, high-quality set of robotic trajectories, we plan to investigate in the future whether incorporating more complex backgrounds can enhance the model’s manipulation performance, either through data generation or further collection efforts. Additionally, although RoboMIND covers a broad range of robot tasks and environments, it currently lacks data from mobile manipulation scenarios. However, since two of our robotic embodiments are mobile, we plan to expand RoboMIND to include mobile manipulation tasks in the future. Additionally, RoboMIND can be further enriched by adding more informative annotations such as high-level planning instructions.
@@ Acknowledgments
We would like to extend our deepest gratitude to the following individuals for their invaluable help in this work: Connor Han, Devin Zhao, Dylan Wang, Emily Chen, Eva Cui, Gasin Wei, Gehrmann Niu, Hailey Huang, Houser Hao, James Zhang, Jack Guo, Jeff Zhang, Jianwei Guo, Jieyu Zhang, Kai Yang, Kora Shu, Magian Wang, Prince Guo, Sheljia Xiao, Shuyi Zhang, Yaowen Xu, Yingjuan Tang, Yizhang Liu, Yoan Zhang, and Zehui Liu. We also sincerely appreciate the dedication and effort of numerous contributors who assisted with data collection, quality assurance, annotation, and testing procedures. Their collective efforts and expertise have been instrumental in making this research possible. We sincerely appreciate their commitment to advancing the field of robotic manipulation through this collaborative endeavor.
@@
  • Initial Drafting: Chengkai Hou, Jiaming Liu, Kun Wu, Meng Li, Yinuo Zhao, Mengzhen Liu, and Xinhua Wang

### v3

@@ VII Limitations
One limitation of RoboMIND is the relatively simple background environments. While our primary focus has been on constructing a large-scale, high-quality set of robotic trajectories, we plan to investigate in the future whether incorporating more complex backgrounds can enhance the model’s manipulation performance, either through data generation or further collection efforts. Additionally, although RoboMIND covers a broad range of robot tasks and environments, it currently lacks data from mobile manipulation scenarios. However, since two of our robotic embodiments are mobile, we plan to expand RoboMIND to include mobile manipulation tasks in the future. Additionally, RoboMIND can be further enriched by adding more informative annotations such as high-level planning instructions and annotations of trajectory quality.
@@ Acknowledgments
We would like to extend our deepest gratitude to the following individuals for their invaluable help in this work: Dapeng Wang, Gang Han, Haifang Huang, Jialiang Shu, Jian Xiao, Jianwei Guo, Jianyu Dong, Jiaxing Wei, Jieyu Zhang, Kai Yang, Kun Niu, Lili Chen, Manzhan Wang, Musen Zhang, Peng Guo, Qiu Cui, Shuyi Zhang, Yaowen Xu, Yijie Guo, Yingjuan Tang, Yizhang Liu, Yue Zhao, Yuheng Zhang, Zehui Liu, Zhen Hao, and Zhennan Zhang. We also sincerely appreciate the dedication and effort of numerous contributors who assisted with data collection, quality assurance, annotation, and testing procedures. Their collective efforts and expertise have been instrumental in making this research possible. We sincerely appreciate their commitment to advancing the field of robotic manipulation through this collaborative endeavor.
@@
  • Paper Writing: Chengkai Hou, Jiaming Liu, Kun Wu, Meng Li, Yinuo Zhao, Mengzhen Liu, Xinhua Wang, Zhengping Che, and Shanghang Zhang